4. Use the following to answer questions a-d:
f(x) = x2 + x and g(x) = x – 5
a. Find h(x) = (f ◦ g)(x).
b. State the domain of h(x) = (f ◦ g)(x).
c. Find h(x) = (g ◦ f)(x).
d. State the domain of h(x) = (g ◦ f)(x).

Answer by user_dude2008(1862)

a) h(x) = (f ◦ g)(x) = f(g(x)) = f(x-5) = (x-5)^2+(x-5) = (x-5)((x-5)+1) = (x-5)(x-4) = x^2 - 9x + 20
Answer: x^2 - 9x + 20

b) Domain: all real numbers

c) h(x) = (g ◦ f)(x) = g(f(x)) = g(x^2+x)= x^2+x-5
Answer: x^2+x-5
d) Domain: all real numbers